OpenStudy (anonymous):

Consider the function f(x) = x - x^2 on the interval (0,1). Which of the following statements is true? a. f is increasing b. f is decreasing c. f is constant d. f is neither increasing, nor decreasing, nor constant

OpenStudy (anonymous):

do you know what the vertex is ?

OpenStudy (anonymous):

Do you know anything about x^2 if 0<x<1? How does it compare to x?

OpenStudy (anonymous):

No that's all the question offers -_-

OpenStudy (anonymous):

actually what i meant was not "did they tell you the vertex" but rather "do you know how to find the vertex?"

OpenStudy (anonymous):

ax^2+bx+c=0 or y=ax^2+bx+ ?

OpenStudy (anonymous):

the vertex of \(y=ax^2+bx+c\) is \(-\frac{b}{2a}\) you have \(y=-x^2+x\) so you your example \(a=-1,b=1\)

OpenStudy (anonymous):

i should be more precise and say "the first coordinate of the vertex is \(-\frac{b}{2a}\)"

OpenStudy (anonymous):

Ok, so where do go from there to get the answer?